OpenStudy (anonymous):

Solve the problem. Instruments on a satellite measure the amount of power generated by the satellite's power supply. The time t and the power P can be modeled by the function P = 50e^{-t/300}, where t is in days and P is in watts. How much power will be available after 378 days? Round to the nearest hundredth. Seleccione una respuesta. a. 1.41 watts b. 141.8 watts c. 14.18 watts d. 0.14 watts

OpenStudy (anonymous):

This a math home work

OpenStudy (anonymous):

means choose one answer

OpenStudy (anonymous):

the answer is50e^(-378/300),use the calculator, you can get

OpenStudy (anonymous):

shd101wyy is correct - the question is asking you to substitute the value for t given in the problem into the equation for P

OpenStudy (anonymous):

but I do not know how to use my calculator LOL, my calculator is not so good

OpenStudy (anonymous):

e^(-378/300)*50 = 14.18

OpenStudy (anonymous):

For the given functions f and g, find the requested composite function value. f(x) = /sqrt{x + 5} , g(x) = 2x ; Find (f ∘ g)(2). Seleccione una respuesta. a. 2 sqrt{7} b. 3 c. sqrt{14} d. 2 sqrt{14}

OpenStudy (anonymous):

is it d right answer?

OpenStudy (anonymous):

What is /sqrt?

OpenStudy (anonymous):

,For the given functions f and g, find the requested composite function value. f(x) = /sqrt{x + 5} , g(x) = 2x ; Find (f ∘ g)(2). Seleccione una respuesta. a. 2 {7} b. 3 c. {14} d. 2{14}

OpenStudy (anonymous):

,For the given functions f and g, find the requested composite function value. f(x) = /{x + 5} , g(x) = 2x ; Find (f ∘ g)(2). Seleccione una respuesta. a. 2 {7} b. 3 c. {14} d. 2{14}